Solve by completing the square:
x^2+6x+4=0
Thanks :)

Answer by stanbon(75887)

Solve by completing the square:
x^2+6x+4=0
-----------
x^2 + 6x + 9 = -4 + 9
(x+3)^2 = 5
---------
x+3 = sqrt(5) or x+3 = -sqrt(5)
x = -3+sqrt(5) or x = -3-sqrt(5)
========================================
Cheers,
Stan H.